OpenStudy (anonymous):

Show all your work. x+y=-4 x-y=2

OpenStudy (lalaly):

x+y=4 ... (1) x-y=2.....(2) from (2) x=2+y substitute in (1) (2+y)+y=4 2y=2 y=1 then substitute in (2) x-1=2 x=3